OpenStudy (webo):

Which of the following is the rational exponent expression of third root of 4n? a. (4n)^3 b. 3n^4 c. (4n)^1/3 d. 4n^1/3

OpenStudy (nick5006):

its C, btw

jhonyy9 (jhonyy9):

what is your opinion please ? do you know the formula sqrt(x) = x^(1/2) so how you see the denominator mean the index of radical

jhonyy9 (jhonyy9):

so using this given formula you see that choice c. is righ,correct sure - exactly how wrote above nick5006

OpenStudy (nick5006):

sorry for leaving without explaining, i was working on a problem of my own. Anyways a radical can be written as a fraction. Taking the rule that \[\sqrt[a]{x^b}=x^{b/a}\]

OpenStudy (welshfella):

Glad you explained - I was going to remind you to do that

OpenStudy (nick5006):

But your trying to find the cube root of 4n, and (4n)^1/3 is the right answer because your taking the cube root of 4n not just "n"... and that's why 4n^1/3 is wrong

OpenStudy (nick5006):

you're*

OpenStudy (nick5006):

@jhonyy9

jhonyy9 (jhonyy9):

yes Nick - i m here